OpenStudy (anonymous):

Please help me! Find the value of 2sqrt(3)/sqrt(10)

jimthompson5910 (jim_thompson5910):

Multiply both top and bottom by \(\Large \sqrt{10}\)

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

20sqrt(30)/sqrt(100) ?

jimthompson5910 (jim_thompson5910):

close, but the sqrt(10) doesn't multiply with the 2 to get 20

jimthompson5910 (jim_thompson5910):

So its 2sqrt(30)/sqrt(100)

jimthompson5910 (jim_thompson5910):

now reduce

jimthompson5910 (jim_thompson5910):

sqrt(100) = ???

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

so 2sqrt(30)/sqrt(100) becomes 2sqrt(30)/10

jimthompson5910 (jim_thompson5910):

2/10 reduces to what???

OpenStudy (anonymous):

1/5

jimthompson5910 (jim_thompson5910):

so the answer is \[\Large \frac{\sqrt{30}}{5}\]

OpenStudy (anonymous):

oh okay, I see where I went wrong. Thanks again

jimthompson5910 (jim_thompson5910):

sure thing